OpenStudy (anonymous):

List the four smallest elements of the set. {y|y = x2 − 1, x is in integers} can some explain to me about this?? I have no idea about this!! I need help I have a quiz tomorrow

OpenStudy (amistre64):

what kind of equation does y measure?

OpenStudy (anonymous):

opss let me rewrite it

OpenStudy (anonymous):

@amistre64 {y|y = x^2 − 1, x is in integers}

OpenStudy (amistre64):

does y=x^2 - 1 form a line, a quadratic, a cubic, is it a sine function? what type of graph do you think we are dealing with?

OpenStudy (anonymous):

I have no idea ... so sorry

OpenStudy (amistre64):

well, then the next best thing to do is to simply make a table of values to work with ....

OpenStudy (amistre64):

you are being asked what the 4 smallest values of y are; given the rule that y=x^2-1 make a table of points to determine how y behaves since youre not able to determine a minimum from recognizing its structure.